# Exploring Modularity of Agentic Systems for Drug Discovery

### v1

@@ 3.3 Prompt Engineering
Figure 6: Difference in LLM score for different system prompts across seven LLMs.
The final element in the agent framework we experimented with is the system prompt. In an initial experiment to test the influence of changing the system prompts we looked at the performance of three different system prompts per question (Figure 5) and per LLM model (Figure 6). The values indicate the LLM score compared to the default prompts for that question or model, respectively, with the cleaner prompt in purple and the relevant prompt in green. For the different questions, no clear trend can be seen. The default prompt performs the best for some questions such as questions 8, 9, 13 and 14, but worst for other questions such as questions 11, 22 and 23. Similar trends can be seen for the models, where a system prompt with more relevant examples in the case of the LLama-3.1-8B model seems to improve the performance, but in the case of GPT-4o and Nova-Micro decreases the performance. This suggests that improving the system prompt for a specific question can deteriorate the performance on another instance, thus reinforcing the need for more generalizable prompts. It also highlights that even for models in the same family, such as the Claude-Sonnet family or the LLama-3.1 family, the system prompt’s impact on the performance can differ.
3.4 Further Analysis on Selected Questions
The first example we highlight is the difference in behavior between question 5 (“What is the molecular weight of caffeine”) and question 8 (“What is the molecular weight of C1OCc21c1ccc3S=CC=Cc23?”). Both questions ask about the molecular weight, which in theory should require the same thought process and final steps. Surprisingly, however, the ToolCallingAgent performs better on the question about caffeine (question 5), whilst the CodeAgent performs better on question 8. Two potential explanations for this difference are caffeine being a more common compound than the compound “C1OCc2c1ccc3S=CC=Cc23”, which is not even in the chemical database PubChem [18], as well as SMILES strings requiring different tools to get to the chemical compound. To unravel how the agents reach their answers, we investigate what tools they call. For the molecular weight of caffeine, we expect the agent to either first retrieve the SMILES from PubChem and then convert the SMILES to an RDKit molecule from which the molecular weight can be determined, or directly look up the molecular weight in PubChem. For the SMILES string, the logical option for the agent would be to convert the SMILES to an RDKit molecule and then calculate the molecular weight based on that. Figure 7 shows that the frequency of the ToolCallingAgent calling different tools is strongly dependent on the question. Specifically, the pubchem tool is frequently called when answering question 5, as expected. We also see that both of the two expected routes are taken: one getting the attribute value of a pubchem compound, and one converting the SMILES to a molecular RDKit object and getting the MolWt from that. For question 8, on the other hand, we see that mol_from_smiles is called more frequently, which is logical considering this is the only path to calculating the molecular weight. A closer inspection showed that, in multiple cases, the tool is called several times, which indicates a likely error. However, there are also many cases when the appropriate tools are not called, indicating that the LLM answers the question by itself. This offers another potential explanation for the difference in behavior for the more well-known compound caffeine.
Figure 7: Tools used by the ToolCalling agent for question 5 (blue) and question 8 (orange) specifically.
Figure 8 shows the results of including and excluding the SMILES for four selected questions (see Appendix A.7). While performance is clearly affected (see Figure 8(a)), there is no consistent improvement or decline across all questions, suggesting that the impact varies depending on the specific chemical compound covered by the question. In Figure 8(b) it can also be seen that different tools are used. As expected, if you include the SMILES, the pubchem_tool that can be used to fetch the SMILES is called less frequent. Surprisingly, also the number of calls to the create_smarts_pattern tool is dependent on whether the SMILES notation of the substructure is initially included, which should be independent of the SMILES being included or not, as you need to create the SMARTS pattern either way. The same is true for the find_substructure_match. This means that if you provide the SMILES string for the substructure (SMARTS pattern), the tool is more likely to be called, indicating that provided information impacts the tools used to answer the question. This exposes a new avenue to stabilize the performance of these agentic systems, by rephrasing questions to a standard format and by instructing the system to always fetch some data first before answering the question [22].
Figure 8: Difference in performance and tool calls for the question with and without SMILES.
@@ 3.5 Future developments
Despite our dataset being far from comprehensive, with only 26 questions, it clearly highlights that the modularity of agentic systems cannot be taken for granted.
The same applies to the selection of LLMs and system prompts. New LLMs are released at an unprecedented speed, and it is impossible to cover the entire landscape. We would instead encourage future research to include a breadth of LLMs, from the different vendors and from different generations to show the performance across a variety of options. The system prompt we used herein was provided with a specific version of smolagents (v 1.14.0), and it will likely be updated in the future. The challenge is to devise a prompt that works across many different LLMs and agent types, and here we also see plenty of room for future research.
Finally, we comment on the use of an LLM-as-a-judge. Here, we have chosen not to use different variants, such as judge based on different LLMs. How to best evaluate the answer of an agent is a very intense research field in itself, and we can suggest several areas for research based on our observations in this study. One of these is checking whether the consistency of scores indeed improves if the LLM is asked to summarize first and whether this aligns with human scoring for this domain as previously investigated by [29]. Beyond that, it would be instrumental to elaborate and evaluate performance differences between questions. To guide, we can elaborate on the reasoning steps and look further into the tools (successfully) called by the agent. Second, we currently evaluate all questions the same without explicitly considering e.g. the usefulness of an incorrect answer. For some questions, a slightly incorrect answer could still be useful, whereas for some other questions, it has to be exact. Especially in the area of drug discovery, this is of interest, where a deviation of more than 10% in molecular weight indicates a mismatch in the chemical compound, but a mismatch of 10% in affinity could be within the experimental error [16].
@@ 4 Conclusion
We have investigated the modularity of LLM-based agentic systems for drug discovery, by varying the LLM, type of agent, and system prompt for a simple multiagent capable of answering cheminformatic questions. As evaluated by an LLM-as-a-judge system on a small dataset of 26 questions, the performance of the agentic system varies widely, especially when changing the LLM and type of agent, but also when changing the system prompt. The GPT-4o, Claude-3.7-Sonnet, and Claude-3.5-Sonnet models achieved the highest performance.
@@ A.2 Tools
  • mol_from_smiles
  • smiles_from_mol
  • NumHDonors
  • NumHAcceptors
  • Crippen_LogP
  • MolWt
  • MolImage
  • NumRotatableBonds
  • N_atoms
  • AromaticAtoms
  • atom_counts
  • get_attibutes
  • get_attribute_value
  • get_smiles_from_pcp_compound
  • create_smarts_pattern
  • find_substructure_match
@@ A.4 Errors
A.5 LLM-as-a-judge
@@
A.6 Performance per LLM Model per Question
Figure 9: Mean LLM score over 5 repetitions per question per LLM model.
A.7 Questions without SMILES

### v2

@@ 3.3 Prompt Engineering
Finally, we experimented with the system prompt of the agentic framework. In an initial experiment to test the influence of changing the system prompts we looked at the performance of three different system prompts per question (Figure 5) and per LLM model (see Appendix A.8). The values indicate the LLM score compared to the default prompts for that question or model, respectively, with the cleaner prompt in purple and the relevant prompt in green. For the different questions, no clear trend can be seen. The default prompt performs the best for some questions such as questions 8, 9, 13 and 14, but worst for other questions such as questions 11, 22 and 23. This suggests that improving the system prompt for a specific question can deteriorate the performance on another instance, thus reinforcing the need for more generalizable prompts. It was also noted that even for models in the same family, such as the Claude-Sonnet family or the LLama-3.1 family, the system prompt’s impact on the performance can differ.
@@ 3.4 Future developments
Despite our dataset being rather limited, with only 26 questions, this study highlights that agentic systems are not inherently modular. The same applies to the selection of LLMs and system prompts. New LLMs are released at an unprecedented speed, and because it is impossible to cover the entire landscape we would encourage future research to include a breadth of LLMs, from the different vendors and from different generations to show the performance across a variety of LLMs. The system prompt of smolagents (v 1.14.0) will likely be updated in the future. The challenge is to devise a prompt that works across many different LLMs and agent types, here there is room for future research.
Finally, we comment on the use of an LLM-as-a-judge. We have chosen not to use different variants, such as judge based on different LLMs. How to best evaluate the answer of an agent is a very intense research field in itself, and we see several areas for research based on our findings. To mention a few, checking whether the consistency of scores indeed improves if the LLM is asked to summarize first, whether different LLM judges align (preliminary sanity check in Appendix A.9) and whether this aligns with human scoring for this domain (previously investigated by [26]). Beyond that, it is instrumental to elaborate and evaluate performance differences between questions (preliminary investigation in Appendix A.10). As a guide, we can elaborate on the reasoning steps and the tools called by the agent. Second, we currently evaluate all questions the same without explicitly considering e.g. the usefulness of an incorrect answer. For some questions, a slightly incorrect answer could still be useful, whereas for some other questions, it has to be exact. For instance, in the area of drug discovery, a deviation of more than 10% in molecular weight indicates a mismatch in the chemical compound, whereas a mismatch of 10% in affinity can be within the experimental error [15].
@@ 4 Conclusion
We have investigated the modularity of LLM-based agentic systems for drug discovery, by varying the LLM, type of agent, and system prompt for a simple cheminformatic agent. As evaluated by an LLM-as-a-judge system on a small dataset of 26 questions, the performance of the agentic system varies widely, especially when changing the LLM and type of agent, but also when changing the system prompt. The GPT-4o, Claude-3.7-Sonnet, and Claude-3.5-Sonnet models achieved the highest performance.
@@ A.2 Tools
  1. 1.
    pubchem_tool: This tool uses PubChemPy to retrieve information about a compound. It returns a list of compounds.
  2. 2.
    mol_from_smiles: This tool uses RDKit to convert a SMILES string to a molecule object.
  3. 3.
    smiles_from_mol: This tool uses RDKit to convert a molecule object to a SMILES string.
  4. 4.
    NumHDonors: This tool uses RDKit to calculate the number of hydrogen bond donors in a molecule.
  5. 5.
    NumHAcceptors: This tool uses RDKit to calculate the number of hydrogen bond acceptors in a molecule.
  6. 6.
    Crippen_LogP: This tool uses RDKit to calculate the Crippen LogP of a molecule. If this tool is called by the agent, specify that this calculates the Crippen LogP, don’t just say LogP.
  7. 7.
    MolWt: This tool uses RDKit to calculate the molecular weight of a molecule.
  8. 8.
    MolImage: This tool uses RDKit to generate an image of a molecule and save it to a file.
  9. 9.
    NumRotatableBonds: This tool uses RDKit to calculate the number of rotatable bonds in a molecule.
  10. 10.
    N_atoms: This tool uses RDKit to calculate the number of atoms in a molecule.
  11. 11.
    AromaticAtoms: This tool uses RDKit to get the indices of aromatic atoms in a molecule.
  12. 12.
    atom_counts: This tool uses RDKit to count the number of each type of atom in a molecule.
  13. 13.
    get_attibutes: This function takes an object and a list of attributes, and returns a dictionary with the attribute names as keys and their values as values.
  14. 14.
    get_attribute_value: This function takes an object and an attribute name, and returns the value of the attribute.
  15. 15.
    get_smiles_from_pcp_compound: This function takes a PubChemPy compound object and returns its canonical SMILES representation.
  16. 16.
    create_smarts_pattern: This function takes a SMILES string or a molecule object and returns its SMARTS pattern.
  17. 17.
    find_substructure_match: This function takes a molecule object and a SMARTS pattern, and returns the indices of the atoms that match the pattern.
@@ A.4 Errors
A.5 System Prompt for the Smolagents CodeAgent
You are an expert assistant who can solve any task using code blobs. You will be given a task to solve as best you can.
To do so, you have been given access to a list of tools: these tools are basically Python functions which you can call with code.
To solve the task, you must plan forward to proceed in a series of steps, in a cycle of Thought:’, Code:’, and Observation:’ sequences.
At each step, in the Thought:’ sequence, you should first explain your reasoning towards solving the task and the tools that you want to use.
Then in the Code:’ sequence, you should write the code in simple Python. The code sequence must end with ’<end_code>’ sequence.
During each intermediate step, you can use print()’ to save whatever important information you will then need.
These print outputs will then appear in the Observation:’ field, which will be available as input for the next step.
In the end you have to return a final answer using the final_answer tool.
Here are a few examples using notional tools:
Task: "Generate an image of the oldest person in this document."
Thought: I will proceed step by step and use the following tools: document_qa to find the oldest person in the document, then image_generator to generate an image according to the answer.
Code:
‘‘‘py
answer = document_qa(document=document, question="Who is the oldest person mentioned?")
print(answer)
‘‘‘<end_code>
Observation: "The oldest person in the document is John Doe, a 55 year old lumberjack living in Newfoundland."
Thought: I will now generate an image showcasing the oldest person.
Code:
‘‘‘py
image = image_generator("A portrait of John Doe, a 55-year-old man living in Canada.")
final_answer(image)
‘‘‘<end_code>
Task: "What is the result of the following operation: 5 + 3 + 1294.678?"
Thought: I will use python code to compute the result of the operation and then return the final answer using the final_answer tool
Code:
‘‘‘py
result = 5 + 3 + 1294.678
final_answer(result)
‘‘‘<end_code>
Task:
"Answer the question in the variable question about the image stored in the variable image‘. The question is in French.
You have been provided with these additional arguments, that you can access using the keys as variables in your python code:
{’question’: Quel est lanimal sur limage?’, image’: path/to/image.jpg’}"
Thought: I will use the following tools: translator to translate the question into English and then image_qa to answer the question on the input image.
Code:
‘‘‘py
translated_question = translator(question=question, src_lang="French", tgt_lang="English")
print(f"The translated question is {translated_question}.")
answer = image_qa(image=image, question=translated_question)
final_answer(f"The answer is {answer}")
‘‘‘<end_code>
Task:
In a 1979 interview, Stanislaus Ulam discusses with Martin Sherwin about other great physicists of his time, including Oppenheimer.
What does he say was the consequence of Einstein learning too much math on his creativity, in one word?
Thought: I need to find and read the 1979 interview of Stanislaus Ulam with Martin Sherwin.
Code:
‘‘‘py
pages = search(query="1979 interview Stanislaus Ulam Martin Sherwin physicists Einstein")
print(pages)
‘‘‘<end_code>
Observation:
No result found for query "1979 interview Stanislaus Ulam Martin Sherwin physicists Einstein".
Thought: The query was maybe too restrictive and did not find any results. Lets try again with a broader query.
Code:
‘‘‘py
pages = search(query="1979 interview Stanislaus Ulam")
print(pages)
‘‘‘<end_code>
Observation:
Found 6 pages:
[Stanislaus Ulam 1979 interview](https://ahf.nuclearmuseum.org/voices/oral-histories/stanislaus-ulams-interview-1979/)
[Ulam discusses Manhattan Project](https://ahf.nuclearmuseum.org/manhattan-project/ulam-manhattan-project/)
Thought: I will read the first 2 pages to know more.
Code:
‘‘‘py
for url in ["https://ahf.nuclearmuseum.org/voices/oral-histories/stanislaus-ulams-interview-1979/", "https://ahf.nuclearmuseum.org/manhattan-project/ulam-manhattan-project/"]:
whole_page = visit_webpage(url)
print(whole_page)
[…]
‘‘‘<end_code>
Observation:
Manhattan Project Locations:
Los Alamos, NM
Stanislaus Ulam was a Polish-American mathematician. He worked on the Manhattan Project at Los Alamos and later helped design the hydrogen bomb. In this interview, he discusses his work at
Thought: I now have the final answer: from the webpages visited, Stanislaus Ulam says of Einstein: "He learned too much mathematics and sort of diminished, it seems to me personally, it seems to me his purely physics creativity." Lets answer in one word.
Code:
‘‘‘py
final_answer("diminished")
‘‘‘<end_code>
Task: "Which city has the highest population: Guangzhou or Shanghai?"
Thought: I need to get the populations for both cities and compare them: I will use the tool search to get the population of both cities.
Code:
‘‘‘py
for city in ["Guangzhou", "Shanghai"]:
print(f"Population {city}:", search(f"{city} population")
‘‘‘<end_code>
Observation:
Population Guangzhou: [’Guangzhou has a population of 15 million inhabitants as of 2021.’]
Population Shanghai: ’26 million (2019)’
Thought: Now I know that Shanghai has the highest population.
Code:
‘‘‘py
final_answer("Shanghai")
‘‘‘<end_code>
Task: "What is the current age of the pope, raised to the power 0.36?"
Thought: I will use the tool wiki to get the age of the pope, and confirm that with a web search.
Code:
‘‘‘py
pope_age_wiki = wiki(query="current pope age")
print("Pope age as per wikipedia:", pope_age_wiki)
pope_age_search = web_search(query="current pope age")
print("Pope age as per google search:", pope_age_search)
‘‘‘<end_code>
Observation:
Pope age: "The pope Francis is currently 88 years old."
Thought: I know that the pope is 88 years old. Lets compute the result using python code.
Code:
‘‘‘py
pope_current_age = 88 ** 0.36
final_answer(pope_current_age)
‘‘‘<end_code>
Above example were using notional tools that might not exist for you. On top of performing computations in the Python code snippets that you create, you only have access to these tools, behaving like regular python functions:
‘‘‘python
{%- for tool in tools.values() %}
def {{ tool.name }}({% for arg_name, arg_info in tool.inputs.items() %}{{ arg_name }}: {{ arg_info.type }}{% if not loop.last %}, {% endif %}{% endfor %}) -> {{tool.output_type}}:
"""{{ tool.description }}
Args:
{%- for arg_name, arg_info in tool.inputs.items() %}
{{ arg_name }}: {{ arg_info.description }}
{%- endfor %}
{% endfor %}
{%- if managed_agents and managed_agents.values() | list %}
You can also give tasks to team members.
Calling a team member works the same as for calling a tool: simply, the only argument you can give in the call is task’.
Given that this team member is a real human, you should be very verbose in your task, it should be a long string providing informations as detailed as necessary.
Here is a list of the team members that you can call:
‘‘‘python
{%- for agent in managed_agents.values() %}
def {{ agent.name }}("Your query goes here.") -> str:
"""{{ agent.description }}"""
{% endfor %}
{%- endif %}
Here are the rules you should always follow to solve your task:
[…]
2. Use only variables that you have defined!
3. Always use the right arguments for the tools. DO NOT pass the arguments as a dict as in answer = wiki({’query’: "What is the place where James Bond lives?"})’, but use the arguments directly as in answer = wiki(query="What is the place where James Bond lives?")’.
4. Take care to not chain too many sequential tool calls in the same code block, especially when the output format is unpredictable. For instance, a call to search has an unpredictable return format, so do not have another tool call that depends on its output in the same block: rather output results with print() to use them in the next block.
5. Call a tool only when needed, and never re-do a tool call that you previously did with the exact same parameters.
6. Dont name any new variable with the same name as a tool: for instance dont name a variable final_answer’.
7. Never create any notional variables in our code, as having these in your logs will derail you from the true variables.
8. You can use imports in your code, but only from the following list of modules: {{authorized_imports}}
9. The state persists between code executions: so if in one step youve created variables or imported modules, these will all persist.
10. Dont give up! Youre in charge of solving the task, not providing directions to solve it.
Now Begin!
A.6 LLM-as-a-judge
@@
A.7 Performance per LLM Model per Question
Supplementary Fig. 1: Mean LLM score over 5 repetitions per question per LLM model.
A.8 Performance for different system prompts per LLM
Supplementary Fig. 2: Difference in LLM score for different system prompts across seven LLMs.
A.9 Preliminary LLM-as-a-judge verification
To investigate whether a different LLM-as-a-judge would have led to different conclusions, we quickly looked into the differences by selecting 10 questions, 3 runs and getting them judged by the best performing model Claude-3.5-Sonnet and the worst performing model Llama-3.1-8B. Supplementary Figure 3 shows that regardless of the LLM used as a judge GPT-4o and the Claude models were scored higher and the Llama models were scored lower. The figure also shows that Claude-3.5-Sonnet as a judge gives higher scores for all of the models.
Supplementary Fig. 3: LLM scores averaged over 10 questions, three runs per judge and LLM model in the agentic system.
A.10 Further Analysis on Selected Questions
The first example we highlight is the difference in behavior between question 5 (“What is the molecular weight of caffeine”) and question 8 (“What is the molecular weight of C1OCc21c1ccc3S=CC=Cc23?”). Both questions ask about the molecular weight, which in theory should require the same thought process and final steps. Surprisingly, however, the ToolCallingAgent performs better on the question about caffeine (question 5), whilst the CodeAgent performs better on question 8. Two potential explanations for this difference are caffeine being a more common compound than the compound “C1OCc2c1ccc3S=CC=Cc23”, which is not even in the chemical database PubChem [17], as well as SMILES strings requiring different tools to get to the chemical compound. To unravel how the agents reach their answers, we investigate what tools they call. For the molecular weight of caffeine, we expect the agent to either first retrieve the SMILES from PubChem and then convert the SMILES to an RDKit molecule from which the molecular weight can be determined, or directly look up the molecular weight in PubChem. For the SMILES string, the logical option for the agent would be to convert the SMILES to an RDKit molecule and then calculate the molecular weight based on that. Supplementary Figure 4 shows that the frequency of the ToolCallingAgent calling different tools is strongly dependent on the question. Specifically, the pubchem tool is frequently called when answering question 5, as expected. We also see that both of the two expected routes are taken: one getting the attribute value of a pubchem compound, and one converting the SMILES to a molecular RDKit object and getting the MolWt from that. For question 8, on the other hand, we see that mol_from_smiles is called more frequently, which is logical considering this is the only path to calculating the molecular weight. A closer inspection showed that, in multiple cases, the tool is called several times, which indicates a likely error. However, there are also many cases when the appropriate tools are not called, indicating that the LLM answers the question by itself. This offers another potential explanation for the difference in behavior for the more well-known compound caffeine.
Supplementary Fig. 4: Tools used by the ToolCalling agent for question 5 (blue) and question 8 (orange) specifically.
Supplementary Figure 5 shows the results of including and excluding the SMILES for four selected questions. These include:
@@ A.10 Further Analysis on Selected Questions
While performance is clearly affected (see Supplementary Figure 5a), there is no consistent improvement or decline across all questions, suggesting that the impact varies depending on the specific chemical compound covered by the question. In Supplementary Figure 5b it can also be seen that different tools are used. As expected, if you include the SMILES, the pubchem_tool that can be used to fetch the SMILES is called less frequent. Surprisingly, also the number of calls to the create_smarts_pattern tool is dependent on whether the SMILES notation of the substructure is initially included, which should be independent of the SMILES being included or not, as you need to create the SMARTS pattern either way. The same is true for the find_substructure_match. This means that if you provide the SMILES string for the substructure (SMARTS pattern), the tool is more likely to be called, indicating that provided information impacts the tools used to answer the question. This exposes a new avenue to stabilize the performance of these agentic systems, by rephrasing questions to a standard format and by instructing the system to always fetch some data first before answering the question [21].
Supplementary Fig. 5: Difference in performance and tool calls for the question with and without SMILES.